3. The probability that an automobile being filled with gasoline also needs an oil change is 0.25; the probability that it needs a new oil filter is 0.40; and the probability that both the oil and the filter need changing is 0.14

(a) If a new oil filter is needed, what is the probability that the oil has to be changed
(b) If the oil has to be changed, what is the probability that a new oil filter is needed?
1
Expert's answer
2020-11-03T16:48:47-0500

Let

probability of needing an oil be O

probability of needing an oil filter be F

"P(O)=0.25"

"P(F)=0.40"

"P(O\cap F)=0.14"

a. "P(O|F)"

"P(O|F)=\frac{P(O\cap F)}{P(F)}"

"=\frac{0.14}{0.4}=0.35"

b. "P(F|O)"

"P(F|O)=\frac{P(O\cap F)}{P(O)}"

"=\frac{0.14}{0.25}=0.56"